El Puerto de Liverpool (Liverpool) is a large retailer in Mexico. The following accounts are selected from its annual report for the fiscal year ended December 31, 2017. For the fiscal year ended December 31, 2017, Liverpool purchased merchandise inventory costing 78,023,979 thousand Mexican pesos. Assume that all purchases were made on account. The following T-accounts reflect information contained in the company’s 2017 and 2016 balance sheets (in thousands of Mexican pesos).

Inventories
12/31/2016 Bal. 13,849,931
12/31/2017 Bal. 18,486,423
Suppliers (Accounts Payable)
15,210,743 12/31/2016 Bal.
22,535,802 12/31/2017 Bal.

a. Prepare the entry, using the financial statement effects template and in journal entry form, to record Liverpool’s purchases for the 2017 fiscal year.

b. What amount did Liverpool pay in cash to its suppliers for the fiscal year ended December 31, 2017? Assume that Suppliers (Accounts payable) is affected only by transactions related to inventory. Answer

c. Prepare the entry, using the financial statement effects template and in journal entry form, to record cost of goods sold for the year ended December 31, 2017.
Note: Use a negative sign with your answers, when appropriate.

Noncash Contra Contrib. Earned
Transaction Cash Asset + Assets - Assets = Liabilities + Capital + Capital Revenues - Expenses = Net income
a. To record purchases for 2017.
c. To record cost of goods sold.
General Journal
Description Debit Credit
a.
To record purchases for 2017.
c.
To record cost of goods sold.
Answer #1
Transaction Cash Asset Noncash Assets Contra Assets = Liabilities + Contrib. Capital Earned Capital Revenues - Expenses = Net
Income
a. To record purchases for 2017. 78023979 78023979
c. To record cost of goods sold. -73387487 -73387487 73387487 -73387487
General Journal
Description Debit Credit
a. Inventories 78023979
Suppliers (Accounts payable) 78023979
To record purchases for 2017.
c. Cost of goods sold 73387487
Inventories 73387487
To record cost of goods sold.

Working:

Cost of goods sold = Inventories beginning balance + Purchases - Inventories ending balance = 13849931 + 78023979 - 18486423 = 73387487

b. Cash paid to suppliers = Suppliers beginning balance + Purchases - Suppliers ending balance = 15210743 + 78023979 - 22535802 = 70698920